OpenStudy (anonymous):

ADD. 3/4 plus 3 5/6??

OpenStudy (across):

3/4 = 9/12 5/6 = 10/12 9/12+10/12=19/12

OpenStudy (chriss):

get an lcd (12) \[\frac{9}{12}+3\frac{10}{12} = 3\frac{19}{12}\] then reduce \[\frac{19}{12} \to 1\frac{7}{12}\] and you have \[4\frac{7}{12}\]

OpenStudy (anonymous):

Thank you!

OpenStudy (chriss):

you are welcome